OpenStudy (anonymous):

Write the equation of the line which passes through (1, –1) and is perpendicular to the line with equation 4y–x=1

OpenStudy (ranga):

4y–x=1 4y = x + 1 y = 1/4x + 1/4 slope of the line = 1/4 slope of perpendicular line is the negative reciprocal of this slope. m = -4 Equation of a line in slope-intercept form is: y = mx + b y = -4x + b It passes therough the point (1,-1) put x = 1 and y = -1 in y = -4x + b and solve for b.

OpenStudy (anonymous):

I figured out the slope (m) is \[\frac{ 1 }{ 4 }\]

OpenStudy (anonymous):

Ok... so, y+1=-4(x-1)

OpenStudy (anonymous):

I got 4x+y=3

OpenStudy (ranga):

Yes. y+1=-4(x-1) is the point-slope form of the line and is correct. They don't tell you which form they want. You can also write the equation in the slope-intercept form in which case you can solve for b as I mentioned above. You can also write the equation in standard form and 4x+y=3 is correct.

OpenStudy (anonymous):

Great! I think they want standard form. Thanks a lot ranga!

OpenStudy (ranga):

You are welcome.